ndretti Company has a single product called a Dak. The company normally produces and sells 86,000 Daks each year at a selling price of $58 per unit. The company’s unit costs at this level of activity are given below:

Direct materials $ 7.50
Direct labor 10.00
Variable manufacturing overhead 2.30
Fixed manufacturing overhead 6.00 ($516,000 total)
Variable selling expenses 4.70
Fixed selling expenses 3.00 ($258,000 total)
Total cost per unit $ 33.50

A number of questions relating to the production and sale of Daks follow. Each question is independent.

Required:

4. Due to a strike in its supplier’s plant, Andretti Company is unable to purchase more material for the production of Daks. The strike is expected to last for two months. Andretti Company has enough material on hand to operate at 25% of normal levels for the two-month period. As an alternative, Andretti could close its plant down entirely for the two months. If the plant were closed, fixed manufacturing overhead costs would continue at 30% of their normal level during the two-month period and the fixed selling expenses would be reduced by 20% during the two-month period.

a. How much total contribution margin will Andretti forgo if it closes the plant for two months?

b. How much total fixed cost will the company avoid if it closes the plant for two months?

c. What is the financial advantage (disadvantage) of closing the plant for the two-month period?

d. Should Andretti close the plant for two months?

5. An outside manufacturer has offered to produce 86,000 Daks and ship them directly to Andretti’s customers. If Andretti Company accepts this offer, the facilities that it uses to produce Daks would be idle; however, fixed manufacturing overhead costs would be reduced by 30%. Because the outside manufacturer would pay for all shipping costs, the variable selling expenses would be only two-thirds of their present amount. What is Andretti’s avoidable cost per unit that it should compare to the price quoted by the outside manufacturer
Answer #1

Calculation of contribution margin per Dak

Selling price

58

Less: variable costs

Direct Material

7.5

Direct Labor

10

Variable Overhead

2.30

Variable Selling Expenses

4.70

Total Variable cost

24.5

Contribution Margin per Unit

33.5

Operating level = 86,000*25%*2/12 = 3583.33 units

4-a. Contribution margin foregone = 3583.33*33.5 = $120,041.56

4-b Fixed cost avoided = 516,000*70%*2/12 + 258,000*20%*2/12 = $68,800

c.Advantage of closing = 68,800-120,041.56 = $(51,241.56) i.e. disadvantage

d.No, should not be closed

5.Calculation of avoidable cost

Direct Material

7.5

Direct Labor

10

Variable Overhead

2.30

Avoidable Fixed manufacturing overhead

1.8

Variable selling expenses avoided

1.57

Avoidable cost per unit

23.17